DOSBox 0.74, Cpu speed: 3000 cycles, Frameskip 0, program: SPECTOC QUIZ: UNKNOWN #7 Molecular Formula: C?H15NO 9 3 3 4000 3000 2000 1000心cm1 ば -1 9 8 76 5 4 3 2 1 0 6 C-Carbon Menu N-Nitrogen Menu 0-0xygen S-Sulfur Menu x-Halogen Menu F1-Help F3-Clr Ans F5-Undo ESC-Quit Menu Press C, N, 0, S, or X
Answer #1

C7H15NO

Double bond equivalence = C + 1 – H/2 + N/2 = 7 + 1 -15/2 + ½ = 1

There is only one unsaturation or double bond in the molecule.

Further from a broad peak between 3200 to 3300 cm-1 and strong peak at around 1000 cm-1 indicates the presence of OH and C-O. Further, there is a C=N at around 1600 cm-1.

InNMR there are 4 types of proton 9 protons showing singlet 1ppm, 3 showing triplet at 1.2 ppm, 3 for a broad singlet and a quartet between 3.7to 4 ppm.

Consistent with the given value the most possible structure is that of